OpenStudy (anonymous):

What are the first three terms of the sequence: a1 = 3 and an = 2(an -1)^2? a. 2, 8, 18 b. 3, 18, 648 c. 3, 32, 50 d. 2, 6, 12, 24

Parth (parthkohli):

First one would be 3. 2 options already eliminated.

OpenStudy (anonymous):

Yeah that's what I thought... I just don't know if the answer is b or c.... Can you please help me?

Parth (parthkohli):

And then a2 would be 2(3)^2 = 2*9 = 18

Parth (parthkohli):

Only one option is left now :D

OpenStudy (anonymous):

So b is the answer? and how'd you get that equation??

Parth (parthkohli):

The equation is given in the question already

OpenStudy (anonymous):

@ParthKohli ARe you really good at math or something?

Parth (parthkohli):

Yes I am :P

Parth (parthkohli):

a2 = 2(a[2 - 1])^2 a2 = 2(a1)^2 a2 = 2(3)^2 a2 = 2 * 9 a2 = 18 That was easy lolol

OpenStudy (anonymous):

haha cool! If i need more help, can i just message you? haha what grade are you in?

Parth (parthkohli):

I'm 12 years old...in 8th grade

OpenStudy (anonymous):

oh wow! so you're like really smart?

OpenStudy (anonymous):

what math are you in??

Parth (parthkohli):

I guess..and that's what everyone says :P

Parth (parthkohli):

I can help you with everything but trig or calculus

OpenStudy (anonymous):

awesome! Ima fan you... i might need more help in a bit hahaa :) thanks!

Parth (parthkohli):

Hmm sure!